# Lifelong Sequential Knowledge Editing without Model Degradation

### v1

@@ Abstract
Prior work in parameter-modifying knowledge editing has shown that large-scale sequential editing leads to significant model degradation. In this paper, we study the reasons behind this and scale sequential knowledge editing to 10,000 sequential edits, while maintaining the downstream performance of the original model. We first show that locate-then-edit knowledge editing methods lead to overfitting on the edited facts. We also show that continuous knowledge editing using these methods leads to disproportionate growth in the norm of the edited matrix. We then provide a crucial insight into the inner workings of locate-then-edit methods. We show that norm-growth is a hidden trick employed by these methods that gives larger importance to the output activations produced from the edited layers. With this “importance hacking”, the edited layers provide a much larger contributions to the model’s output. To mitigate these issues, we present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing. ENCORE controls for overfitting and the disproportionate norm-growth to enable long-term sequential editing, where we are able to perform up to 10,000 sequential edits without loss of downstream performance. ENCORE is also 61% faster than MEMIT and 64% faster than AlphaEdit on Llama3-8B.
Machine Learning, ICML
@@ 1 Introduction
(a) MEMIT
(b) AlphaEdit
@@
Figure 1: The continuous growth of norm of edited MLP matrices in LLama3-8B during sequential knowledge editing, as a function of number edits.
Knowledge editing is the task of editing specific facts that a model learns during pretraining in a data and computing efficient manner (De Cao et al., 2021; Yao et al., 2023; Hartvigsen et al., 2024). In this paper, we focus on a category of parameter-modifying knowledge editing methods called “locate-then-edit” methods, which modify only a small subset of the model parameters to add or update knowledge. These methods have been the focus of many recent works (Gu et al., 2024; Gupta et al., 2024a, c; Ma et al., 2024; Kolbeinsson et al., 2024; Fang et al., 2024). While prior work showed knowledge editing methods like ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) lead to a catastophic loss of downstream performance within a few hundred sequential edits (Gu et al., 2024; Gupta et al., 2024b), a recently introduced method called AlphaEdit (Fang et al., 2024) has been able to push this to 3000 edits. In this paper, we aim to enable large-scale sequential knowledge editing without causing model degradation. Towards this goal, we push sequential knowledge editing to what we propose as the next frontier - performing 10,000 sequential knowledge edits without loss of downstream performance.
Our work is based on two main observations. Firstly, we show that existing locate-then-edit knowledge editing methods are prone to overfitting on the edited fact, where the output probability of an edited fact is unusually higher when compared to the confidence with which an unedited model predicts facts. Secondly, we show that sequential knowledge updates made to a model consistently lead to an increase in the norm of the weight matrix being edited, as shown in Figure 1. We show that this norm-growth is a secret trick used by locate-then-edit methods, which we call “importance hacking”. The increasing norm of the edited matrix also increases the norm of the activation vectors produced by the edited layers on average. This allows the outputs produced from the edited layers to have a larger influence on the final output of the model. With this, the edited layers are able to override the information coming from other parts of the model, leading to successful knowledge edits but inadvertently causing a loss of general ability, which might require information coming from other parts of the model.
Based on these observations, we present targeted interventions to overcome the limitations of overfitting on edited facts and disproportionate norm growth. We first formalize locate-then-edit methods as a two-step finetuning process, where the first step uses gradient descent for optimization and the second step directly optimizes the loss using a least-square objective. This way of understanding knowledge editing enable us to apply appropriate interventions depending on the optimization scenario. To mitigate overfitting during the gradient descent process, we propose Most-Probable Early Stopping (MPES) - where we halt gradient descent when edited facts become most probable across all the different contexts used to calculate the loss. MPES reduces overfitting on edited facts and consequently improves editing performance while reducing loss of downstream performance. To counteract the growth of the norm of the edited matrix during sequential knowledge editing, we add a frobenius-norm constraint within the knowledge editing objective of locate-then-edit methods (Meng et al., 2022b). This constraint still allows for a closed-form solution and gives us the benefits of computationally efficient knowledge edits while allowing us to control the increasing norm of edited weight matrices, thus enabling longer and more robust sequential knowledge editing.
Towards large-scale sequential knowledge editing, we present ENCORE - Early stopping and Norm-Constrained Robust knowledge Editing. ENCORE combines MPES with Frobenius-norm constrained objective and enables performing 10,000 sequential knowledge edits while maintaining the downstream performance of the original unedited model. ENCORE significantly outperforms prior locate-then-edit methods like ROME, MEMIT, and AlphaEdit, as verified on GPT2-XL, Llama-2-7B, and Llama-3-8B, and is significantly faster than its predecessors. Specifically, ENCORE is 61% faster than MEMIT and 64% faster than AlphaEdit when editing Llama3-8B.
2 Background and Related Work
In this section, we provide a brief introduction to locate-then-edit knowledge editing methods and present them as a two-step fine-tuning process. For a more detailed introduction to these methods, we refer the reader to prior works (Meng et al., 2022a, b; Gupta et al., 2024c).
Locate-then-edit family of methods like ROME (Meng et al., 2022a) , MEMIT (Meng et al., 2022b) and AlphaEdit (Fang et al., 2024) are used to update facts that can be represented in the form of triplets of the form (subject, relation, object) or $(s,r,o)$. Instead of updating all the weights of a model to incorporate new knowledge, these methods only update certain weight matrices that are most responsible for factual recall (Meng et al., 2022a). The location of an edit within a model is described by a two-dimensional address - (i) an intermediate layer to be edited and (ii) a token from the list of input tokens used to create the target representation. The exact layer to be edited is found using causal tracing (Meng et al., 2022a) or an empirical sweep over all decoder layers of the model (Hase et al., 2024; Yoon et al., 2024). Additionally, updating the second MLP layer in the FFN module of decoder layers has shown optimal knowledge editing performance (Geva et al., 2020; Meng et al., 2022a). This provides the first part of the editing address. Meng et al. (2022a) also showed that using the output representation of the subject token of a sentence produces the best editing results. This provides the second part of the editing address, where the edit is made using the position index of the last token of the subject. We explain this with an example below.
@@ 2 Background and Related Work
(a) Gradient descent step which finds the target activations for the MLP matrix.
@@
Given a fact to be edited, for example - “The capital of Malaysia is Singapore”, the query phrase for the editing process is “The capital of Malaysia is” and the target phrase is “Singapore”. The first part of the editing address, the exact layer whose second MLP matrix gets edited, is decided before the editing begins. The second part of the editing address is the token index of the last subject token, which in this case would be the last subword-token in “Malasiya”. The intermediate hidden representation of this last subject token is used to make the edit.
Once the editing address has been decided, instead of updating the chosen MLP weight matrix directly using gradient descent, the process of locate-then-edit knowledge editing proceeds in two steps -
@@
    In the first step (Figure 2(a)), gradient descent is used to find the appropriate activation vector that acts as a target for the weight matrix to be edited. This target activation is found such that the edited fact is generated by the model. In the example, this activation will cause the model to generate “Singapore” in response to the question. Note that in this step, no weights are updated and just an intermediate activation vector is found. This new activation vector now serves as the target for the MLP weight matrix chosen for editing.
@@
    The weight update happens in the second step of editing (Figure 2(b)), where the MLP matrix is updated with the target activation vector found in the previous step, using a least squares loss function. This loss function tries to preserve the outputs of the MLP matrix for unrelated contexts while generating the target activation when the input corresponds to the query phrase.
Specifically, let $W_{0}$ be the initial weights of the second MLP matrix, which is being edited to $\hat{W}$. $k_{0}$ is used to indicate an input to the MLP matrix representing activation vectors for information we want to preserve from the original model, and $k_{e}$ is input activation vectors representing facts we want to insert into the model. Let $v_{e}$ be the desired target activation vector for the edited MLP matrix found in step 1 of editing using gradient descent. Then the loss function used to update the MLP weight matrix is formulated using least-squares in the form of a preservation-memorization objective (Gupta et al., 2024c):
@@
Since the above objective is linear in the argument, we do not need to use gradient descent for optimization. Thus, locate-then-edit methods can be seen as a unique type of fine-tuning method. Instead of updating the MLP matrix directly using gradient descent on the desired data, the weight update happens in two steps using two different types of objective functions for each step. The first step uses gradient descent whereas the second step uses a closed-form solution.
3 Methods, Models, Datasets and Evaluation
In this paper, we focus on three “locate-then-edit” methods - ROME (Meng et al., 2022a), MEMIT (Meng et al., 2022b) and AlphaEdit (Fang et al., 2024). As discussed in section 2, each of these algorithms use gradient descent to find intermediate target representations followed by a second optimization step which is linear in the argument. ROME uses an equality-constraint to enforce memorization in place of the least square constraint as shown in equation 1, whereas equation 1 represents the objective function used in MEMIT. AlphaEdit adds a null-space projection term in the MEMIT objective.
Prior work has shown model degradation during sequential editing with ROME and MEMIT (Gu et al., 2024; Gupta et al., 2024b), whereas AlphaEdit is a recent method that is able to perform sequential editing for up to 3,000 facts. In this paper, we adopt the experimental setting of AlphaEdit where they perform sequential edits in batches of 100 facts. This means that 100 facts are edited into the model with each weight update, and multiple such updates are performed sequentially. Since ROME only allows for one edit to be made at a time to the model, we used the batched generalization of ROME, called EMMET (Gupta et al., 2024c), which uses the same equality-constraint objective as ROME but generalizes it to the batched editing case.
We evaluate all algorithms on three representative models - GPT2-XL (Radford et al., 2019), Llama2-7B (Touvron et al., 2023) and Llama3-8B (Yoon et al., 2024). All experiments are performed on the CounterFact (Meng et al., 2022a) and zsRE (Mitchell et al., 2021) datasets, which are standard knowledge editing datasets. We present the results for Llama2-7B and Llama3-8B on CounterFact dataset in the main paper and present the remaining results in the appendix due to space constraints.
In this paper, we evaluate the editing algorithms along two dimensions - editing performance and downstream performance. The editing performance evaluates the success of the knowledge editing algorithm in making successful edits, while downstream performance evaluates the extent of model degradation following prior work (Fang et al., 2024; Gupta et al., 2024b; Gu et al., 2024).
Table 1: Comparison between prediction probabilities of edited facts versus facts that a model knows through pretraining. MPES results in more natural prediction probabilities while being 39% - 76% faster than other methods.
@@ 3 Methods, Models, Datasets and Evaluation
Method Model Original Fact Prob Edited Fact Prob w/o MPES Edited Fact Prob w/ MPES Time Per Edit w/o MPES (s) Time Per Edit w/ MPES (s)
EMMET GPT2-XL 0.39 0.98 0.47 1.26 0.63 ($\downarrow$ 50%)
Llama2-7B 0.52 0.92 0.59 6.61 1.86 ($\downarrow$ 71%)
Llama3-8B 0.49 0.99 0.61 7.65 1.86 ($\downarrow$ 76%)
MEMIT GPT2-XL 0.39 0.65 0.14 1.60 0.97 ($\downarrow$ 39%)
Llama2-7B 0.52 0.75 0.41 4.84 2.79 ($\downarrow$ 42%)
Llama3-8B 0.49 0.77 0.41 8.71 3.31 ($\downarrow$ 61%)
AlphaEdit GPT2-XL 0.39 0.98 0.39 1.49 0.83 ($\downarrow$ 44%)
Llama2-7B 0.52 0.84 0.37 5.89 2.69 ($\downarrow$ 54%)
Llama3-8B 0.49 0.75 0.33 9.44 3.43 ($\downarrow$ 63%)
Knowledge Editing Metrics:
To evaluate editing performance, we use five standard knowledge editing metrics (Meng et al., 2022a). (i) Efficacy Score (ES), which measures if an edit was successfully made, (ii) Paraphrase Score (PS), which measures if the model is able to recall edited facts in different scenarios, (iii) Neighborhood Score (NS), which measures if edited facts disturbs unrelated knowledge, (iv) Overall Score (S), which is the harmonic mean of ES, PS and NS, and (v) Generation Entropy (GE), which measures the fluency of a model. A detailed explanation of these metrics is given in Appendix A.
The editing metrics for each model are calculated after making all 10,000 sequential edits (Fang et al., 2024). This approach ensures that the metrics capture both the success of the edits of the latest batch of facts as well as facts that were edited in the past.
Downstream Performance Metrics:
Following prior work (Fang et al., 2024; Gupta et al., 2024b), we measure downstream performance during knowledge editing using 6 tasks - massive multitask language understanding (MMLU) (Hendrycks et al., 2020), natural language inference (NLI, RTE) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007), sentiment analysis (SST2) (Socher et al., 2013), paraphrase detection (MRPC) (Dolan & Brockett, 2005), and linguistic acceptability classification (CoLA) (Warstadt et al., 2019). We measure the downstream performance of models every 1000 edits, following the experimental settings of Fang et al. (2024). More details about experimental details can be found in Appendix J.
4 Overfitting during Knowledge Editing

### v2

# Lifelong Knowledge Editing requires Better Regularization

@@ Abstract
Knowledge editing is a promising way to improve factuality in large language models, but recent studies have shown significant model degradation during sequential editing. In this paper, we formalize the popular locate-then-edit methods as a two-step fine-tuning process, allowing us to precisely identify the root cause of this degradation. We show that model degradation occurs due to (1) over-optimization of internal activations and (2) continuous norm-growth of edited matrices. To mitigate these issues, we introduce two regularization techniques: (1) Most-Probable Early Stopping (MPES) and (2) explicit Frobenius norm-constraint. We demonstrate that applying these simple yet effective regularization techniques at key points in the editing process can substantially mitigate model degradation. Combining these regularization methods enables scaling locate-then-edit methods to 10,000 edits while reducing editing time by 42-61%. These results show that targeted regularization is essential for lifelong knowledge editing.
Lifelong Knowledge Editing requires Better Regularization
Akshat Gupta1*, Phudish Preteepamornkul1,2*, Maochuan Lu1 Ahmed Alaa1, Thomas Hartvigsen3, Gopala Anumanchipalli1 1UC Berkeley, 2SCB DataX, 3University of Virginia akshat.gupta@berkeley.edu
@@ 1 Introduction
Knowledge editing entails editing specific facts that a language model has learned during pre-training in a data and computate efficient manner De Cao et al. (2021); Yao et al. (2023). The most popular editors follow a “locate-then-edit” approach, where the methods aim to locate and update small subsets of a model’s parameters that recall target facts Meng et al. (2022a, b); Gupta et al. (2024a, c); Ma et al. (2024); Fang et al. (2024). Although these methods have shown strong results when performing singular or small-scale edits, they suffer substantial degradation when applied at scale Gu et al. (2024a); Gupta et al. (2024b); Thede et al. (2025), leaving the problem of lifelong knowledge editing largely unsolved. * Equal Contribution
(a) Norm-growth as function of number of edits
(b) Layer wise norm post-editing
@@
Figure 1: The continuous growth of norm of edited MLP matrices in MEMIT Llama3-8B during sequential knowledge editing.
In this work, we systematically investigate the factors that cause model degradations and propose appropriate regularization methods to mitigate them, enabling large-scale sequential knowledge editing without sacrificing downstream performance. We begin by first formalizing locate-then-edit methods as a two-step fine-tuning process (section 4), where the first step uses gradient-descent for activation optimization, and the second step performs weight-update via a least-squares update rule. This explicit structure allows us to precisely diagnose the causes of loss of downstream performance and allows us to choose appropriate interventions.
We show that model degradation occurs during continuous sequential editing because of two reasons. Firstly, we show that the gradient-descent step in locate-then-edit methods leads to over-optimization of internal activations, which makes the model predict edited facts with unnaturally high confidence (section 5). We mitigate this by proposing Most-Probable Early Stopping (MPES) - a novel variant of early-stopping where we halt gradient-descent steps when edited facts become most probable across all the different contexts used to calculate the loss. Secondly, we show that sequential editing leads to disproportionate growth of the Frobenius norm of the edited matrix (Figure 1). This allows the outputs produced from the edited layers to have an abnormally larger influence on the final output of the model, inadvertently causing loss of general ability, which might require information coming from other parts of the model. To address this, we incorporate a Frobenius norm-constraint into the editing objective of locate-then-edit methods (section 6.2). While early stopping and norm-constraints are known regularization methods, their targeted stage-specific application to locate-then-edit methods is non-trivial. With our work, we demonstrate that there exists a lack of proper regularization in existing knowledge editing methods and show that explicit regularization at the appropriate stages of editing is essential and critical for scalable and stable knowledge editing. The code for our work can be found here - https://github.com/scalable-model-editing/knowledge-editing-regularization
Our proposed regularization methods mitigate the respective issues of over-optimization of intermediate activation vectors and disproportionate norm-growth, consequently preserving downstream performance for a larger number of edits. Finally, we show that combining these two regularization methods allows us to perform up to 10,000 sequential edits while maintaining original downstream performance and making editing 41-62% faster.
2 Related Work
Knowledge editing methods can broadly be classified into three types (Yao et al., 2023; Zhang et al., 2024c) - hypernetwork based editing methods (De Cao et al., 2021; Mitchell et al., 2021), where an additional model is trained to predict the edited weights of the model, in-context editing methods (Mitchell et al., 2022; Hartvigsen et al., 2024; Zhong et al., 2023), where edited information is added in-context, and locate-then-edit methods Meng et al. (2022a, b); Gupta et al. (2024c, a); Fang et al. (2024), where the knowledge source is traced to certain MLP layers Hase et al. (2024); Yoon et al. (2024) before modifying them using a two-step fine-tuning method (section 4). This work focuses on performing locate-then-edit type of knowledge editing methods.
Overfitting During Knowledge Editing.
Some recent works have studied overfitting as a key challenge of knowledge editing. For example, Zhang et al. (2024a) identifies that locate-then-edit methods overfit on edited facts. They propose to mitigate this by introducing an inference-stage regularization approach called “learn to inference” (LTI) that guides edited models to recall new knowledge by adding multi-stage constraints during optimization. With our proposed variant of early-stopping called "most-probable early stopping" (MPES), we present a much simpler way of reducing overfitting. We also show that LTI does not scale to multiple edits and leads to immediate model degradation when performing sequential editing. Meanwhile, MPES leads to significant improvements in downstream performance while countering overfitting (section 5).
Norm-Growth and Weight Regularization.
Prior work has shown that sequential knowledge editing leads to the norm-growth of the updated weight matrices Gupta et al. (2024b); Yang et al. (2024). Ma et al. (2025) present a similar observation in the form of growth of condition number of the edited matrices as a potential cause of loss of downstream performance. However, none of these works explains why increasing norm of edited matrix is bad for the model. In our work, we analyze the effect of norm-growth on internal activations and show how it connects with model degradations (section 6.1). We verify our conclusions by adding a norm-constraint in the optimization objective of MEMIT. We show that our proposed norm-constraint combines seamlessly with MEMIT’s closed-form solution, unlike prior regularization methods like PRUNE Ma et al. (2025) and RECT Gu et al. (2024b), which require additional regularization steps post-editing. We also show that our methods significantly outperform PRUNE and RECT in preserving downstream performance over a large number of edits.
3 Methods, Models, and Evaluation
We adopt the experimental setting of AlphaEdit Fang et al. (2024), a recent locate-then-edit method that is able to perform sequential editing for up to 3,000 facts. Following their setting, we perform sequential edits in batches of 100 facts. This means that 100 facts are edited into the model with each weight update, and multiple such updates are performed sequentially.
We evaluate all algorithms on three representative models - GPT2-XL Radford et al. (2019), Llama2-7B Touvron et al. (2023) and Llama3-8B Yoon et al. (2024). All experiments are performed on the CounterFact Meng et al. (2022a) and zsRE Mitchell et al. (2021) datasets, which are standard knowledge editing datasets. We present the results for Llama2-7B and Llama3-8B on the CounterFact dataset in the main paper and present the remaining results in the appendix due to space constraints.
In this paper, we evaluate the editing algorithms along two dimensions - editing performance and downstream performance. The editing performance evaluates the success of the knowledge editing algorithm in making successful edits, while downstream performance evaluates the extent of model degradation following prior work Fang et al. (2024); Gupta et al. (2024b); Gu et al. (2024a).
@@ 3 Methods, Models, and Evaluation
(a) gradient-descent step which finds the target activations for the MLP matrix.
@@
Knowledge Editing Metrics:
To evaluate editing performance, we use five standard knowledge editing metrics Meng et al. (2022a). (i) Efficacy Score (ES), which measures if an edit was successfully made, (ii) Paraphrase Score (PS), which measures if the model is able to recall edited facts in different scenarios, (iii) Neighborhood Score (NS), which measures if edited facts disturb unrelated knowledge, (iv) Overall Score (S), which is the harmonic mean of ES, PS, and NS, and (v) Generation Entropy (GE), which measures the fluency of a model. A detailed explanation of these metrics is given in Appendix A. The editing metrics for each model are calculated after making all 10,000 sequential edits Fang et al. (2024). This approach ensures that the metrics capture both the success of the edits of the latest batch of facts as well as facts that were edited in the past.
Downstream Performance Metrics:
Following prior work Fang et al. (2024); Gupta et al. (2024b), we measure downstream performance during knowledge editing using 6 tasks - massive multitask language understanding (MMLU) Hendrycks et al. (2020), natural language inference (NLI, RTE) Dagan et al. (2005); Haim et al. (2006); Giampiccolo et al. (2007), sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), and linguistic acceptability classification (CoLA) Warstadt et al. (2019). Performance is assessed every 1000 edits, following Fang et al. (2024). Additional details are provided in Appendix N.
4 Knowledge Editing as a Two-Step Fine-tuning Process
“Locate-then-edit" family of methods like ROME Meng et al. (2022a) , MEMIT Meng et al. (2022b) and AlphaEdit Fang et al. (2024) update factual knowledge expressed as (subject, relation, object) or (s,r,o). Instead of updating all parameters of a model to incorporate new knowledge, these methods only update certain matrices that are most responsible for fact recall Meng et al. (2022a). The location of an edit within a model is described by a two-dimensional address - (i) an intermediate layer to be edited and (ii) a token from the list of input tokens used to create the target representation.
The exact layer to be edited is found via causal tracing Meng et al. (2022a) or an empirical sweep across the model layers Hase et al. (2024); Yoon et al. (2024). Only the second MLP matrix in the FFN module of the editing layer is updated (Geva et al., 2020; Meng et al., 2022a), forming the first part of the editing address. Meng et al. (2022a) also showed that using the output representation of the last subject token produces the best editing results, providing the second part of the address. The entire editing process is presented in Figure 2.
We explain this with an example. Given a fact to be edited, for example - “The capital of Malaysia is Singapore", the query phrase for the editing process is “The capital of Malaysia is" and the target phrase is “Singapore". The first part of the editing address, the exact layer whose second MLP matrix gets edited, is decided before the editing begins. The second part of the editing address is the token index of the last subject token, which in this case would be the last subword-token in “Malasiya". The intermediate hidden representation of this last subject token is used to make the edit. Once the editing address has been decided, instead of updating the chosen MLP weight matrix directly using gradient-descent, locate-then-edit knowledge editing proceeds in two steps.
@@ 4 Knowledge Editing as a Two-Step Fine-tuning Process
    In the first step (Figure 2(a)), gradient-descent is used to find the appropriate activation vector that acts as a target for the weight matrix to be edited. In the example, the found activation will cause the model to generate “Singapore" in response to the question. The loss function for the gradient-descent step is shown in equation 2. Note that in this step, no weights are updated and just an intermediate activation vector is found.
@@
    The weight update occurs in the second editing step (Figure 2(b)), where the MLP matrix is updated using the target activation vector found previously. This update uses a least square loss function, which preserves the MLP outputs for unrelated contexts while generating the target activation when the input corresponds to the query phrase.
The loss function used to update the MLP weight matrix is formulated using least-squares in the form of a preservation-memorization objective Gupta et al. (2024c):
@@
Specifically, $W_{0}$ represents the initial weights of the second MLP matrix, which is being edited to $\hat{W}$. $k_{0}$ represent input to the MLP matrix for information we want to preserve from the original model, and $k_{e}$ is input activation vectors representing facts we want to insert into the model. $v_{e}$ is the target activation vector found in step 1 of editing using gradient-descent. Since the above objective is linear in the argument, we do not need to use gradient-descent for optimization.
@@
Thus, locate-then-edit methods can be seen as a unique type of 2-step fine-tuning method. Instead of updating the MLP matrix directly using gradient-descent on the desired data, the weight update happens in two steps using two different types of objective functions for each step. The first step uses gradient-descent whereas the second step uses a closed-form solution.
@@
Model Prediction Probability Time / edit (s) Original fact MEMIT MEMIT w/ LTI MEMIT w/ MPES MEMIT MEMIT w/ LTI MEMIT w/ MPES Llama-2 7B 0.52 0.78 0.30 0.45 4.84 8.67 2.79 ($\downarrow$ 42 %) Llama-3 8B 0.49 0.79 0.29 0.41 8.71 9.21 3.31 ($\downarrow$ 61 %)
Table 1: Comparison between prediction probabilities of edited facts along with editing time when using MPES and LTI Zhang et al. (2024b) with MEMIT.
5 Over-Optimization of Target Activations